OpenStudy (anonymous):

Find the sum of the series. 1/2^n. n=1 (summation symbol) to infinity.

OpenStudy (anonymous):

i know its a geometric series but am not sure how they found a= 1/2 and r= 1/2. I know the formula for a geometric series is a/(1-r)....and that r has to be less than 1.

OpenStudy (amistre64):

1/2 + 1/2^2 + 1/2^3 + ... + 1/2^n 1/2 (1 + 1/2 + 1/2^2 + ... + 1/2^n-1) \[1/2\frac{1-1/2^n}{1-1/2}\] since |r|<1 r^inf = 0 \[1/2\frac{1}{1-1/2}\]

OpenStudy (anonymous):

oh ok, thanks!